OpenStudy (anonymous):

Which of the following is in simplest form? twenty-two thirty-fifths forty-two fifty-fourths nine forty-fifths twenty-four twenty-sevenths

OpenStudy (kagıtucak):

you always ask nearly the same type of questions :)

OpenStudy (anonymous):

.....

OpenStudy (kagıtucak):

twenty-four twenty-sevenths

OpenStudy (kagıtucak):

because in the first one30/5=620 +6=26 simplest second one 20/4=5 40 +5=45 simplest third one 40/5=8 9+8=17 simplest

OpenStudy (anonymous):

thx ^_^